The price of a computer is decreasing at an exponential rate. In 1985, a desktop computer cost $2,500. The price of a comparable computer is decreasing by 3.5% every 2 years. During what year is the price of that comparable computer $1,200?
Answer by stanbon(75887)

The price of a computer is decreasing at an exponential rate. In 1985, a desktop computer cost $2,500. The price of a comparable computer is decreasing by 3.5% every 2 years. During what year is the price of that comparable computer $1,200?
------
1200 = 2500(1-0.035)^(t/2)
(12/25) = (0.0965)^(t/2)
-----
Take the log of both sides::
(t/2)*log(0.0965) = log(12/25)
t/2 = -0.3188/-0.0155 = 20.6
time = 41.2 years
Ans:: 1985+42 = 2027
-----------
Cheers,
Stan H.